The formula r = 2(sqrt of 5L) can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid going 50 miles per hour?
Answer by RAY100(1637)

r= 2*sqrt(5L)
.
for r=50
.
50 =2*sqrt(5L)
.
25= sqrt(5L),,,,square both sides
.
625 = 5L
.
125 ft = L
.
check,,,r= 2*sqrt(5*125) = 2*sqrt625 = 2*25 =50,,,,,ok
.